The sum of 2 numbers is 98. The difference of the same 2 numbers is 40. What are the 2 numbers?
Answer by macston(5194)

.
x and y are two numbers
x-y=40
x=40+y Use this to substitute for x
x+y=98 Substitute for x from above.
(40+y)+y=98 Subtract 40 from each side
2y=58
y=29 ANSWER 1: one of the numbers is 29.
x=40+y=29+y=69 Answer 2: the other number is 62.
CHECK:
x-y=40
69-29=40
40=40
x+y=98
69+29=98
98=98